# NeurDB: An AI-powered Autonomous Data System

### v1

@@ Abstract
In the wake of rapid advancements in artificial intelligence (AI), we stand on the brink of a transformative leap in data systems. The imminent fusion of AI and DB (AI×$\times$×DB) promises a new generation of data systems, which will relieve the burden on end-users across all industry sectors by featuring AI-enhanced functionalities, such as personalized and automated in-database AI-powered analytics, self-driving capabilities for improved system performance, etc. In this paper, we explore the evolution of data systems with a focus on deepening the fusion of AI and DB. We present NeurDB, our next-generation data system designed to fully embrace AI design in each major system component and provide in-database AI-powered analytics. We outline the conceptual and architectural overview of NeurDB, discuss its design choices and key components, and report its current development and future plan.
@@ 1 Introduction
In this position paper, we trace the evolutionary trajectory of data systems and characterize related works on AI×$\times$×DB, establishing the timeliness and necessity of advancing toward next-generation data systems. We showcase promising AI-powered applications for which we believe the next-generation data system has much to offer. Through a comprehensive analysis of user needs derived from these applications, we identify key challenges in deepening the fusion of AI and DB to realize next-generation data systems. Firstly, enabling in-database AI-powered analytics to render data systems inherently AI-friendly proves to be intricate, particularly in accommodating modern AI applications like Large Language Models (LLMs) and generative AI. Secondly, the vision of fully autonomous, intelligent data systems is still a distant reality, as existing data systems necessitate extensive manual oversight and maintenance. Lastly, the fusion of AI and databases introduces new challenges in data privacy, which existing techniques often overlook.
We present NeurDB, our initiative towards creating a next-generation system that seamlessly supports AI applications and fully embraces AI techniques. We introduce methodologies and insights to address the above challenges, focusing on creating an in-database AI-friendly ecosystem, developing an intelligent self-driving data system, and providing privacy preservation and trusted AI×$\times$×DB. For each objective, we delve into various new techniques employed in NeurDB, including in-database dynamic model selection and slicing, fast-adaptive learned database components, in-database federated learning, etc. We also outline future pathways for the development of NeurDB, highlighting the open research challenges that must be addressed along this journey.
@@ 2.2 Related work on AI×$\times$×DB
For one thing, databases can significantly enhance AI applications. Databases have the potential to efficiently manage numerous AI models using their inherent user-friendly SQL interfaces, and provide high-quality data with their transaction guarantees to enhance AI model accuracy. For another, AI can substantially optimize databases. AI-driven techniques, such as knob tuning, can automatically refine database configurations to improve performance, which often surpasses traditional manual tuning done by database administrators (DBAs). However, existing approaches on AI×$\times$×DB typically focus on optimizing within their specific domains, and have not yet embraced a full fusion of AI and DB into a unified data system. We now delve deeper into the existing works to highlight the integration challenges and opportunities.
1) DB support for AI. Database techniques under this category aim to improve data quality for more accurate model training, provide automatic model selection, and accelerate model inference. However, existing approaches are often built on top of the database rather than fully integrated within it, which limits their effectiveness and efficiency. First, while various approaches, such as data cleaning, data integration, data discovery, data enrichment, and data lineage, are proposed to improve data quality, they are not directly supported by existing databases. Implementing these techniques often requires complex combinations of multiple SQL commands, which can be user-unfriendly and inefficient. Second, several systems [5] that are built on top of databases can automatically select suitable models for AI applications. However, these systems often need to retrieve and transfer data from the database, which can be costly and time-consuming due to the increasing volume of data. Third, recent proposals [6, 7] target to enhance model inference efficiency, however, they typically assume static models that can be buffered in memory, which is not always the case with today’s dynamic, large-scale models.
2) AI support for DB. AI has shown potential for optimizing database systems. Previous works [8] introduce automatic knob tuning to improve system performance, but they provide limited insight into the possibilities of AI-powered database components, leaving broader opportunities for AI enhancements largely untapped. Despite several efforts to create learned database components, such as learned query optimizers for more accurate query plans and learned indexes for faster key searches, these solutions rely heavily on offline training on static datasets. This paradigm limits their effectiveness when dealing with the dynamic workloads common in modern applications, rendering them impractical for real-world use. Further, existing works often neglect the potential of combining AI with recent advancements in databases, such as the disaggregated architecture that separates computation from storage in cloud environments to improve scalability and elasticity.
@@ 2.3 Emerging AI-powered applications
AI-powered applications are increasingly becoming integral across various sectors, which aim to relieve the burden on end-users by offering personalized, automated AI-powered analytics and predictions. Here we outline the transformative impact of AI in healthcare and e-commerce domains.
@@ 2.3.1 Intelligent healthcare
Healthcare has continuously improved over the years, and with recent advancements in AI techniques, it is becoming smarter. We envision that the future of AI-powered healthcare promises exceptional intelligence, enabling more precise diagnoses, customized treatment plans, and proactive health management [9] AI agents will play a crucial role in assisting physicians in achieving more precise diagnoses. By leveraging AI models trained on extensive datasets, including genetic profiles, medical imaging, electronic health records, physicians’ notes, etc., these agents can reveal patterns not detectable by the human eye in areas such as neurology and orthopedics. Further, AI-powered virtual assistants will offer 24/7 support to patients. They will handle inquiries about medication schedules, offer advice on managing chronic conditions, provide insights into physiological changes, and ensure adherence to treatment plans.
@@ 3.1 Key design principles
Data systems are dynamic in nature as the data distributions and query patterns evolve over time, and therefore, the systems have to adapt for the dynamicity while guaranteeing reliability.
@@
Due to the unpredictable nature of real-world environments and the inherent uncertainties in prediction accuracy and latency, the data systems must adapt and provide acceptable performance. The system optimization workflow often incurs several iterations. We are therefore using the Filter-and-Refine Principle (FRP) to guide the design of NeurDB. For a dynamic system with huge amounts of data, FRP is the basis for adaptivity, efficiency, and scalability. The filter-and-refine strategy is typically used to speed up computationally intensive tasks so that they can be completed within the pre-defined latency and/or resource constraints. In some tasks, such as a join, we have to remove data that are surely not contributing to the answer quickly, and then perform computationally intensive steps on a much smaller subset of the original data. Likewise, we eliminate query processing plans that are obviously suboptimal quickly, and only look for a near-optimal plan. For iterative computations, we should take actions that move to the target within a smaller search space.
@@ 3.2 Key features
Next-generation data systems are confronted with three critical challenges that must be addressed to efficiently and effectively support AI-powered applications.
@@ 3.2.2 Intelligent self-driving data system
We envision that in-database AI-powered analytics will catalyze a comprehensive evolution of data system architecture and techniques. This transformation requires AI optimizations to make the data system scalable, adaptive, and intelligent.
@@ 3.3 Conceptual framework of NeurDB
To address these challenges, we establish the conceptual framework of NeurDB in Figure 3, which underpins its overall design and functionality.
@@ 4.3 Dynamic model slicing
Model slicing is a critical technique for adjusting AI models to meet the specific requirements of various analytical tasks. It customizes the architecture of deep learning models based on analytical queries, thus improving the model inference accuracy. As data volumes grow, analytical queries often focus on specific subsets of the data rather than the entire database. Traditionally, a single general model is used to support inference for different analytical queries [43, 44]. However, such a model, optimized to capture common patterns across the entire database, may not deliver as accurate inferences as a dedicated model trained on a specific subset of interest. For example, when examining patient data, analysts prefer to study patterns or trends among individuals diagnosed with a particular disease. A model exclusively trained from the data of diabetic patients might provide more precise predictions than the general model for all patients regarding their readmission rates. Given the vast number of possible analytical queries, training a dedicated model for each one is also computationally impractical. To this end, we consider leveraging query conditions to perform model slicing. In particular, for each analytical query, we analyze and encode the query conditions, which indicate the specific data required for the query, and then use this information to adjust the model architecture accordingly.
@@ 5 AI-powered autonomous data system
In this section, we present the architecture of autonomous NeurDB, and detail the key components that we have developed to transform NeurDB into an AI-powered autonomous data system.
@@ 5.1 Autonomous architecture overview
Developing such an efficient and effective autonomous data system is not trivial, particularly given the dynamic nature of workloads. The equipped AI model must adapt to these variations to maintain accuracy while minimizing impact on system performance. For example, while learning from specific instances of datasets and query patterns can potentially obtain improved query plans, such improvements may be short-lived due to workload shifts and data updates. We therefore need to design new techniques to overcome these challenges. In particular, we aim to design dynamic and online AI models that facilitate continuous optimization of system components for better performance. Given that the Filter-and-Refine paradigm (FRP) is used as a common-known strategy in indexing [53, 54, 55], query processing [56, 57], etc., our core idea generally follows the FRP and employs online reinforcement learning (RL) to facilitate dynamicity. RL naturally aligns with FRP, as it involves filtering through a range of possible actions and then refining the most effective strategies based on feedback. In particular, FRP’s filtering stage is akin to RL’s exploration phase, where the RL agent tries different actions to understand their effects on the environment, gathering information to guide future decisions. Further, the exploitation phase of RL shares the same objectives as the refinement step of the FRP, where the RL agent refines its strategy by focusing on actions that yield higher rewards. During this phase, the RL agent also optimizes its strategy to maximize long-term gains by learning from past experiences. With these design principles in mind, we now elaborate on our key techniques for the query optimizer and concurrency control to realize seamless online adaptation in NeurDB.

### v2

@@ Abstract
In the wake of rapid advancements in artificial intelligence (AI), we stand on the brink of a transformative leap in data systems. The imminent fusion of AI and DB (AI×$\times$×DB) promises a new generation of data systems, which will relieve the burden on end-users across all industry sectors by featuring AI-enhanced functionalities, such as personalized and automated in-database AI-powered analytics, self-driving capabilities for improved system performance, etc. In this paper, we explore the evolution of data systems with a focus on deepening the fusion of AI and DB. We present NeurDB, an AI-powered autonomous data system designed to fully embrace AI design in each major system component and provide in-database AI-powered analytics. We outline the conceptual and architectural overview of NeurDB, discuss its design choices and key components, and report its current development and future plan.
@@ 1 Introduction
In this position paper, we trace the evolutionary trajectory of data systems and characterize related works on AI×$\times$×DB, establishing the timeliness and necessity of advancing toward next-generation AI-powered autonomous data systems. We showcase promising AI-powered applications for which we believe the next-generation data system has much to offer. Through a comprehensive analysis of user needs derived from these applications, we identify key challenges in deepening the fusion of AI and DB to realize next-generation data systems. Firstly, enabling in-database AI-powered analytics to render data systems inherently AI-friendly proves to be intricate, particularly in accommodating modern AI applications like Large Language Models (LLMs) and generative AI. Secondly, the vision of fully autonomous, intelligent data systems is still a distant reality, as existing data systems necessitate extensive manual oversight and maintenance. Lastly, the fusion of AI and databases introduces new challenges in data privacy, which existing techniques often overlook.
We present NeurDB, our initiative towards creating an AI-powered autonomous system that provides in-database AI-powered analytics to seamlessly support AI applications, and fully embraces AI techniques in each major system component to offer self-driving capabilities. We introduce methodologies and insights to address the above challenges, focusing on creating an in-database AI-friendly ecosystem, developing an intelligent self-driving data system, and providing privacy preservation and trusted AI×$\times$×DB. For each objective, we delve into various new techniques employed in NeurDB, including in-database dynamic model selection and slicing, fast-adaptive learned database components, in-database federated learning, etc. We also outline future pathways for the development of NeurDB, highlighting the open research challenges that must be addressed along this journey.
@@ 2.2 Related work on AI×$\times$×DB
For one thing, databases can significantly enhance AI applications. Databases have the potential to efficiently manage numerous AI models using their inherent user-friendly SQL interfaces [5], and efficiently manage training/validation data for AI models [1]. For another, AI can substantially optimize databases. AI-driven techniques, such as knob tuning, can automatically refine database configurations to improve performance, which often surpasses traditional manual tuning done by database administrators (DBAs). However, existing approaches on AI×$\times$×DB typically focus on optimizing within their specific domains, and have not yet embraced a full fusion of AI and DB into a unified data system. We now delve deeper into the existing works to highlight the integration challenges and opportunities.
1) DB support for AI. Database techniques under this category aim to improve data quality for more accurate model training, provide automatic model selection, and facilitate model inference. However, existing approaches are often built on top of the database rather than fully integrated within it, which limits their effectiveness and efficiency. First, while various approaches, such as data cleaning, data integration, data discovery, data enrichment, and data lineage, are proposed to improve data quality, they are not directly supported by existing databases. Implementing these techniques often requires complex combinations of multiple SQL commands, which can be user-unfriendly and inefficient. Second, several systems [5, 6] that are built on top of databases can automatically select suitable models for AI applications. However, these systems often need to retrieve and transfer data from the database, which can be costly and time-consuming due to the increasing volume of data. Third, recent proposals [7, 8] target to enhance model inference efficiency, however, they typically focus on traditional machine learning models, which may be ineffective for today’s large-scale deep learning models.
2) AI support for DB. AI has shown potential for optimizing database systems. Previous works [9, 10] introduce automatic knob tuning to improve system performance, but they provide limited insight into the possibilities of AI-powered database components, leaving broader opportunities for AI enhancements largely untapped. Despite several efforts to create learned database components, such as learned query optimizers for more accurate query plans and learned indexes for faster key searches, these solutions rely heavily on offline training on static datasets. This paradigm limits their effectiveness when dealing with dynamic workloads in real applications, rendering them impractical for real-world use. Further, existing works often neglect the potential of combining AI with recent advancements in databases, such as the disaggregated architecture that separates computation from storage in cloud environments to improve scalability and elasticity.
@@ 2.3 Emerging AI-powered applications
AI-powered applications are increasingly becoming integral across various sectors, which aim to relieve the burden on end-users by offering personalized, automated AI-powered analytics and predictions. With NeurDB, the construction of such AI applications can be simplified, and their efficiency can be improved. Here we outline the transformative AI applications in healthcare and e-commerce domains, and envision how they can be built with the unique functionalities of NeurDB.
@@ 2.3.1 Intelligent healthcare
Healthcare has continuously improved over the years, and with recent advancements in AI techniques, it is becoming smarter. We envision that the future of AI-powered healthcare promises exceptional intelligence, enabling more precise diagnoses, customized treatment plans, and proactive health management [11]. AI-powered healthcare assistants will play a crucial role in assisting physicians in achieving more precise diagnoses. By leveraging AI models trained on extensive datasets, including genetic profiles, medical imaging, electronic medical records (EMRs), physicians’ notes, etc., these assistants can reveal patterns not detectable by the human eye in areas such as neurology and orthopedics. More importantly, these virtual assistants can offer 24/7 support to patients. They will handle inquiries about medication schedules, offer advice on managing chronic conditions, provide insights into physiological changes, and ensure adherence to treatment plans.
NeurDB can simplify the establishment of intelligent healthcare applications. Specifically, the digital diagnosis system, where doctors and physicians record EMR data and other relevant patient and treatment information, can be constructed based on NeurDB, which provides a mature infrastructure for fast data access and processing. Since all the required data is stored in NeurDB, applications such as AI-powered healthcare assistants can directly perform the necessary and predictive analysis within NeurDB using its in-database AI-powered analytic capabilities, which will be elaborated in Section 3. In general, future intelligent healthcare applications can be efficiently managed by NeurDB, eliminating the complexity of involving multiple dedicated systems for data management, model serving, etc.
@@ 2.3.2 Intelligent e-commerce
We envision that NeurDB can efficiently support intelligent e-commerce applications. For example, NeurDB can support online purchases by maintaining complex relational tables of purchasing records and context attributes. Based on such relational tables, we can conduct predictive tasks, such as link prediction to indicate potential purchases and fraud detection, directly within NeurDB. In particular, NeurDB can automatically transform the relational purchasing table into a graph with user and item nodes, and then build deep learning models, such as GNNs or transformers, on this graph. This approach simplifies the process compared to traditional methods where graph construction and model development are performed outside the database, often resulting in higher costs and complexity. Moreover, fine-tuning a pre-trained LLM based on the underlying domain-specific e-commerce data is also desirable and feasible within NeurDB.
@@ 3.1 Key design principles
Data systems are dynamic in nature as the data distributions and query patterns evolve over time, and therefore, the systems have to adapt for the dynamicity while guaranteeing reliability. We are therefore using the Filter-and-Refine Principle (FRP) to guide the design of NeurDB.
@@
Due to the unpredictable nature of real-world environments and the inherent uncertainties in prediction accuracy and latency, the data systems must adapt and provide acceptable performance. The system optimization workflow often incurs several iterations. For a dynamic system with huge amounts of data, FRP is the basis for adaptivity, efficiency, and scalability. The filter-and-refine strategy is typically used to speed up computationally intensive tasks so that they can be completed within the pre-defined latency and/or resource constraints. In some tasks, such as a join, we have to remove data that are surely not contributing to the answer quickly, and then perform computationally intensive steps on a much smaller subset of the original data. Likewise, we eliminate query processing plans that are obviously suboptimal quickly, and only look for a near-optimal plan. For iterative computations, we should take actions that move to the target within a smaller search space.
@@ 3.2 Key features
Next-generation data systems are confronted with three critical features that must be achieved to efficiently and effectively support AI-powered applications.
@@ 3.2.2 Intelligent self-driving data system
We envision that in-database AI-powered analytics will catalyze a comprehensive evolution of data system architecture and techniques, leading to a fully autonomous data system that can dynamically adapt itself with little or even without human intervention. This transformation requires AI optimizations to make the data system scalable, adaptive, and intelligent.
@@ 3.3 Conceptual framework of NeurDB
To provide these key features, we establish the conceptual framework of NeurDB in Figure 3, which underpins its overall design and functionality. NeurDB consists of four key components: ① in-database AI-friendly ecosystem, ② intelligent self-driving data system, ③ privacy-preserving and trusted AI×$\times$×DB, and ④ automated performance modeling. These components correspond with and interact with each other to effectively and efficiently support ⑤ AI-powered applications in various domains.
@@ 4.3 Dynamic model slicing
Model slicing [44] is a critical technique for adjusting AI models to meet the specific requirements of various analytical tasks. It customizes the architecture of deep learning models based on analytical queries, thus improving the model inference accuracy. As data volumes grow, analytical queries often focus on specific subsets of the data rather than the entire database. Traditionally, a single general model is used to support inference for different analytical queries [45, 46]. However, such a model, optimized to capture common patterns across the entire database, may not deliver as accurate inferences as a dedicated model trained on a specific subset of interest. For example, when examining patient data, analysts prefer to study patterns or trends among individuals diagnosed with a particular disease. A model exclusively trained from the data of diabetic patients might provide more precise predictions than the general model for all patients regarding their readmission rates. Given the vast number of possible analytical queries, training a dedicated model for each one is also computationally impractical. To this end, we consider leveraging query conditions to perform model slicing. In particular, for each analytical query, we analyze and encode the query conditions, which indicate the specific data required for the query, and then use this information to adjust the model architecture accordingly.
@@ 5 AI-powered autonomous data system
In this section, we present the architecture of autonomous NeurDB, and detail the key components that we have developed to transform NeurDB into an intelligent self-driving data system.
@@ 5.1 Autonomous architecture overview
Developing such an efficient and effective autonomous data system is not trivial, particularly given the dynamic nature of workloads. The major challenge is that as data in the database evolves over time, static AI models can become inefficient in providing reliable results. For example, while learning from specific instances of datasets and query patterns can potentially obtain improved query plans, such improvements may be short-lived due to workload shifts and data updates. We therefore aim to design dynamic and online AI models that can be updated and adapted to data variations to maintain accuracy, thereby facilitating continuous optimization of system components for better performance. Given that the Filter-and-Refine paradigm (FRP) is used as a common-known strategy in indexing [55, 56, 57], query processing [58, 59], etc., our core idea generally follows the FRP and employs online reinforcement learning (RL) to facilitate dynamicity. RL naturally aligns with FRP, as it involves filtering through a range of possible actions and then refining the most effective strategies based on feedback. In particular, FRP’s filtering stage is akin to RL’s exploration phase, where the RL agent tries different actions to understand their effects on the environment, gathering information to guide future decisions. Further, the exploitation phase of RL shares the same objectives as the refinement step of the FRP, where the RL agent refines its strategy by focusing on actions that yield higher rewards. During this phase, the RL agent also optimizes its strategy to maximize long-term gains by learning from past experiences. With these design principles in mind, we now elaborate on our key techniques for the query optimizer and concurrency control to realize seamless online adaptation in NeurDB.